A spring has a force constant of B0 N/m, it vibrates with a frequency of 7 Hz. Calculate the,

a) period;

b) angular frequency:

c) mass of the spring,


1
Expert's answer
2021-02-04T10:19:38-0500

Answer

a) period

"T=\frac{1}{T}=\frac{1}{7}=0.14" sec

b) angular frequency

"w=\frac{2\pi}{T}=48.8rad" /sec

c)

mass of the spring,

"m=\frac{k}{w^2}=0.04kg"